# Extending Whisper with Prompt Tuning to Target-Speaker ASR

### 13 Dec 2023 v1

@@ Abstract
Target-speaker automatic speech recognition (ASR) aims to transcribe the desired speech of a target speaker from multi-talker overlapped utterances. Most of the existing target-speaker ASR (TS-ASR) methods involve either training from scratch or fully fine-tuning a pre-trained model, leading to significant training costs and becoming inapplicable to large foundation models. This work leverages prompt tuning, a parameter-efficient fine-tuning approach, to extend Whisper, a large-scale single-talker ASR model, to TS-ASR. Variants of prompt tuning approaches along with their configurations are explored and optimized for TS-ASR. Experimental results show that prompt tuning can achieve performance comparable to state-of-the-art full fine-tuning approaches while only requiring about 1% of task-specific model parameters. Notably, the original Whisper’s features, such as inverse text normalization and timestamp prediction, are retained in target-speaker ASR, keeping the generated transcriptions natural and informative.
@@ 2.1 Background of Whipser
$\begin{split}{\bf H}=&~{}{\rm AudioEncoder}_{\phi_{e}}({\bf X}),\\ \hat{y}_{t}=&~{}{\rm TextDecoder}_{\phi_{d}}(\mathbf{g},\hat{y}_{1:t-1},\mathbf{H}).\end{split}$
@@
2.2 Extending Whipser to Target-speaker ASR
@@ 2.2.1 Prompt Tuning
Prompt tuning was first proposed in NLP to augment a frozen large language model (LLM) for new tasks [21, 22]. This is achieved by introducing a limited number of leading special tokens to the input text. The special tokens function similarly to $\bf g$ in Whipser, which offers the flexibility of turning a foundation model into different operation modes between multiple tasks. Moreover, these tokens are assigned with trainable token embeddings, termed soft prompts [14, 15], which are randomly initialized and optimized by stochastic gradient descent.
Formally, given the log-Mel spectrogram $\mathbf{\widetilde{X}}$ of an overlapped speech from $U$ speakers, we extend Whipser to transcribe the text token sequence $\mathbf{\hat{y}}_{i}$ of the $i$-th speaker, conditioned on the corresponding speaker embedding $e_{i}\in\mathbb{R}^{d_{e}}$:
$\begin{split}\mathbf{H}_{i}=&~{}{\rm AudioEncoder}_{\phi_{e}}({[\mathbf{W}e_{i},\mathbf{P}_{e},\mathbf{\widetilde{X}}]}),\\ \hat{y}_{i,t}=&~{}{\rm TextDecoder}_{\phi_{d}}(\mathbf{g}(\mathbf{P}_{d}),\hat{y}_{i,1:t-1},\mathbf{H}_{i}),\end{split}$ (1)
where $\mathbf{W}\in\mathbb{R}^{d_{m}\times d_{e}}$ is a trainable weight matrix; $\mathbf{P}_{e}\in\mathbb{R}^{d_{m}\times L_{e}}$ and $\mathbf{P}_{d}\in\mathbb{R}^{d_{m}\times L_{d}}$ are learnable soft prompts; $L_{e}$ and $L_{d}$ are hyper-parameters denoting the number of soft prompts; and $\mathbf{g}(\mathbf{P}_{d})$ represents a modification of the special tokens $\mathbf{g}$, i.e., soft prompts $\mathbf{P}_{d}$ are inserted into the prev-text of $\mathbf{g}$.
@@
$(\mathbf{W^{*}},\mathbf{P^{*}})=\operatorname*{arg\,min}_{\mathbf{W},\mathbf{P}}\quad\mathcal{L}_{CE}(\mathbf{\hat{y}}_{i},\mathbf{y}_{i}).$ (2)
@@ 2.2.2 Deep Prompting
In (1), soft prompts are inserted into the encoder and decoder input. We observe that such a simple strategy may have the following limitations: First, it constrains the number of tunable parameters because the prompt length cannot be too long for computational cost concerns. Second, it is difficult to directly control the model prediction since these soft prompts are “far” from the output layer. Therefore, we further explore using deep prompting [20], wherein we replace the preceding output vectors from intermediate blocks with new soft ones for performance improvement; see Fig. 1 for an illustration.
@@ 2.2.3 Reparameterization of Soft Prompts
It has been reported that directly optimizing soft prompts may make the training process unstable, resulting in poor performance [15, 16]. In order to improve the stability of the training process, the reparameterization approach can be adopted [16]. It uses a residual feed-forward network to reparameterize $\mathbf{P}$ as:
@@ 3.2 Training Configuration
We evaluate the performance of our prompting scheme on three models: Whisper-large v2, Whisper-medium, and Whisper-small. The dimensions of prompts are set to be equal to the width of each model, specifically, 1280 for Whisper-large v2, 1024 for Whisper-medium, and 768 for Whisper-small.
In the training phase, we fix all the parameters of Whisper and optimize the soft prompts, the speaker projection layer, and the reparameterization MLP (if it exists) jointly using the AdamW optimizer [28] with an initial learning rate of 1e-4. We train a total of 10 epochs on train-100-clean and another epoch on train-100-both. After the first 5 epochs, we reduce the learning rate to 1e-5. The training process is performed on a small server with one single NVIDIA RTX3090 24GB.
@@ 3.3.1 Selection of Prompt Length and Reparameterization Method
We first evaluate how the prompt length and reparameterization method affect the performance. Specifically, we choose Whisper-large v2 as the foundation model and extend it to TS-ASR using deep prompting while varying prompt lengths within {4, 8, 16, 32, 64}. Additionally, we evaluate the performance of two different reparameterization schemes: one that employs a shared MLP to reparameterize all prompts, and another that uniquely employs separate MLPs for each layer. As a point of reference for benchmarking, we also examine a scheme that does not involve the utilization of reparameterization MLPs.
We compute the average WER for the three reparameterization methods under different prompt lengths. The results are shown in Table 1. It is observed that as the prompt length increases, the WER performance tends to improve. On the other hand, once the length exceeds 16, the performance gains from further length increase become less prominent. Therefore, to strike a balance between performance and parameter efficiency, we opt to set the prompt length to 16 for all subsequent experiments.
@@
Table 1 also shows how the three reparameterization schemes influence the performance. It is observed that employing a shared reparameterization MLP leads to performance degradation even compared to the scheme without reparameterization MLP. This could be attributed to that the shared MLP restricts the learning of inter-layer variations. This is proved by the fact that adopting separate MLPs for each layer leads to performance enhancements. As a result, we opt to employ separate MLPs for the reparameterization of soft prompts in our subsequent experiments.
3.3.2 Performance and Parameter Efficiency Comparison With SOTAs
In this section, we compare our prompting scheme with other SOTA methods for TS-ASR in two aspects: performance (evaluated in terms of WER) and parameter efficiency (evaluated in terms of task parameters). The results are shown in Table 2. It is observed that deep prompting achieves comparable performance with existing methods; it is promising to see that such a performance is achieved by only requiring a very limited amount of task-specific parameters.
In addition, employing reparameterization MLP gives an additional performance improvement on test-clean at the expense of an increase in the number parameters to tune. It is worth mentioning that when the training is completed, the MLP can be immediately discarded. In this way, the number of task-specific parameters to be stored for inference remains small.
@@ 3.3.2 Performance and Parameter Efficiency Comparison With SOTAs
We then report the results of full fine-tuning in the last two rows of Table 2. Due to memory constraints, we are unable to provide the results for Whisper-large. A noticeable trend emerges when comparing full fine-tuning with the prompting approach: despite a substantial increase in the number of tunable parameters, full fine-tuning fails to exhibit satisfactory performance. The underlying cause of this performance gap is attributed to the training set train-100-clean is relatively small, causing severe over-fitting during full fine-tuning. This stark contrast underscores the effectiveness of prompt tuning in mitigating overfitting risks, especially when adaptation data for large-scale models are limited.
@@ 3.3.3 Ablation Study
We conduct ablation experiments to verify the impact of design components, including prompt tuning, deep prompting, and reparameterization MLP. As shown in Table 3, Whisper without any modification (i.e. zero-shot configuration) cannot perform well on target-speaker ASR. When we insert target speaker embedding together with soft prompts into both encoder and decoder input, a dramatic WER reduction, e.g. from 78.92% to 17.59% for Whisper-large on test-clean, is observed. This shows the efficacy of prompt tuning. By further adding the reparameterization MLP module, we observe performance improvement in most of the cases, showing that reparameterization of soft prompts can benefit the training. In addition, configurations with deep prompting show consistent performance improvement. This demonstrates that added tunable parameters can enhance the performance and the soft prompts close to the output layer have a more direct control on model prediction.
We further investigate how prompting location affects performance. Specifically, based on the best configuration (i.e., Whisper + PT + DP + MLP), we ablate the soft prompts inserted in the encoder/decoder to test the influence of varying prompting location. Here, we set the prompt length to 32 to keep the number of parameters constant. The results are shown in the last two rows of Table 3, from which we see that the soft prompts inserted in encoder play an important role to guide the model while those inserted in decoder play a relatively minor role but still lead to performance improvement. So, inserting soft prompts both in encoder and decoder will lead to the best performance with a constant number of parameters.
@@ 3.3.4 Retaining Whisper’s Featured Abilities
We anticipate a practical TS-ASR system to not only accurately transcribe the desired speech of a target speaker but also to execute the entire speech processing pipeline. This includes various post-processing tasks, such as inverse text normalization and timestamp tagging, among others. The original Whisper has already demonstrated commendable performance in these aspects. However, it’s important to note that the prompted Whisper can no longer showcase its automatic punctuating and capitalizing abilities during performing TS-ASR. This limitation arises because the ground truth texts in the LibriSpeech dataset, which are manually labeled, lack case sensitivity and punctuation. Soft prompts have learned and perpetuated this bias. To retain the automatic inverse text normalization ability of Whisper, we propose to employ the auto-labeled texts generated by the original Whisper model on single-talker utterances as ground truth to supervise the training of soft prompts.
@@
The above table shows the results with and without inverse text normalization, where we highlight the automatically inserted punctuation and capitalized letters. We see that through training soft prompts supervised by the auto-labeled texts that are inverse-text-normalized, the prompted Whisper can now automatically perform inverse text normalization along with target-speaker ASR with virtually no performance loss. By further removing the special token $\langle$|⟨ |no-timestamps|$\rangle$| ⟩ in multitask format (see Sec. 2.1), the prompted Whisper can produce reliable timestamps, as indicated by the highlights in the last row of the table, even though timestamps were not included in the adaptation data.
@@ 4 Conclusion
In this work, we have proposed a prompting scheme, which jointly leverage prompt tuning, deep prompting together with reparameterization technique, to extend the single-talker ASR model to the target-speaker ASR of multi-talker overlapped utterances. Our proposed method has shown comparable performance with SOTA and such performance was achieved with an attractively small amount of task parameters compared to existing methods. Notably, our method also retained the automatic inverse text normalization and timestamp prediction capabilities of the original Whisper model. This ensures that the generated transcriptions remain both natural and informative, further enhancing the versatility and utility of our method.

### 11 Jan 2024 v2

@@ Abstract
Target-speaker automatic speech recognition (ASR) aims to transcribe the desired speech of a target speaker from multi-talker overlapped utterances. Most of the existing target-speaker ASR (TS-ASR) methods involve either training from scratch or fully fine-tuning a pre-trained model, leading to significant training costs and becoming inapplicable to large foundation models. This work leverages prompt tuning, a parameter-efficient fine-tuning approach, to extend Whisper, a large-scale single-talker ASR model, to TS-ASR. Variants of prompt tuning approaches along with their configurations are explored and optimized for TS-ASR. Experimental results show that prompt tuning can achieve performance comparable to state-of-the-art full training approaches while only requiring about 1% of task-specific model parameters. Notably, the original Whisper’s features, such as inverse text normalization and timestamp tagging, are retained in target-speaker ASR, keeping the generated transcriptions natural and informative.
@@ 2.1 Background of Whisper
$\begin{split}{\bf H}=&~{}{\rm AudioEncoder}_{\phi_{e}}(Conv({\bf X})),\\ \hat{y}_{t}=&~{}{\rm TextDecoder}_{\phi_{d}}(\mathbf{g},\hat{y}_{1:t-1},\mathbf{H}).\end{split}$
@@
2.2 Extending Whisper to Target-speaker ASR
@@ 2.2.1 Prompt Tuning
Prompt tuning was first proposed in NLP to augment a frozen large language model (LLM) for new tasks [21, 22]. This is achieved by introducing a limited number of leading special tokens to the input text. The special tokens function similarly to $\bf g$ in whisper, which offers the flexibility of turning a foundation model into different operation modes between multiple tasks. Moreover, these tokens are assigned with trainable token embeddings, termed soft prompts [14, 15], which are randomly initialized and optimized by stochastic gradient descent.
Formally, given the log-Mel spectrogram $\mathbf{\widetilde{X}}$ of an overlapped speech from $U$ speakers, we extend whisper to transcribe the text token sequence $\mathbf{\hat{y}}_{i}$ of the $i$-th speaker, conditioned on the corresponding speaker embedding $e_{i}\in\mathbb{R}^{d_{e}}$:
$\begin{split}\mathbf{H}_{i}=&~{}{\rm AudioEncoder}_{\phi_{e}}({[\mathbf{W}e_{i},\mathbf{P}_{e},Conv(\mathbf{\widetilde{X}})]}),\\ \hat{y}_{i,t}=&~{}{\rm TextDecoder}_{\phi_{d}}(\mathbf{g}(\mathbf{P}_{d}),\hat{y}_{i,1:t-1},\mathbf{H}_{i}),\end{split}$ (1)
where $\mathbf{W}\in\mathbb{R}^{d_{m}\times d_{e}}$ is a trainable weight matrix; $\mathbf{P}_{e}\in\mathbb{R}^{d_{m}\times L_{e}}$ and $\mathbf{P}_{d}\in\mathbb{R}^{d_{m}\times L_{d}}$ are learnable soft prompts; $L_{e}$ and $L_{d}$ are hyper-parameters denoting the number of soft prompts; and $\mathbf{g}(\mathbf{P}_{d})$ represents a modification of the special tokens $\mathbf{g}$, i.e., soft prompts $\mathbf{P}_{d}$ are inserted into the token embeddings of the prev-text in $\mathbf{g}$.
@@
$(\mathbf{W^{*}},\mathbf{P^{*}})=\operatorname*{arg\,min}_{\mathbf{W},\mathbf{P}}\quad\mathcal{L}_{\text{C}E}(\mathbf{\hat{y}}_{i},\mathbf{y}_{i}).$ (2)
@@ 2.2.2 Deep Prompting
In Eq. (1), soft prompts are inserted into the encoder and decoder input. We observe that such a simple strategy may have the following limitations: First, it constrains the number of tunable parameters because the prompt length cannot be too long for computational cost concerns. Second, it is difficult to directly control the model prediction since these soft prompts are “far” from the output layer. Therefore, we further explore using deep prompting [20], wherein we replace the preceding output vectors from intermediate blocks with new soft ones for performance improvement; see Fig. 1 for an illustration.
@@ 2.2.3 Reparameterization of Soft Prompts
It has been reported that directly optimizing soft prompts may make the training process unstable, resulting in poor performance [15, 16]. To improve the stability of the training process, the reparameterization approach can be adopted [16]. It uses a residual feed-forward network to reparameterize $\mathbf{P}$ as:
@@ 3.2 Training Configuration
We evaluate the performance of our prompting scheme on three models: Whisper-large v2, Whisper-medium, and Whisper-small.In the training phase, we fix all the parameters of Whisper and optimize the soft prompts, the speaker projection layer, and the reparameterization MLP (if it exists) jointly using the AdamW optimizer [28] with an initial learning rate of 1e-4. We train a total of 10 epochs on train-100-clean and another epoch on train-100-both. After the first 5 epochs, we reduce the learning rate to 1e-5. The training process is performed on a small server with one single NVIDIA RTX3090 24GB. For all experiments, We keep random seeds the same for fair comparison.
@@ 3.3.1 Selection of Prompt Length and Reparameterization Method
We first analyze how the prompt length and reparameterization method affect the performance. We choose Whisper-large v2 as the foundation model and extend it to TS-ASR using deep prompting while varying prompt lengths within {4, 8, 16, 32, 64}. Additionally, we evaluate the performance of two reparameterization schemes: (1) sharedMLP that employs a shared MLP to reparameterize all prompts; and (2) sepMLP that uniquely employs separate MLPs for each layer. As a point of reference for benchmarking, we also examine a scheme that does not involve the utilization of reparameterization MLPs.
@@
The results are shown in Table 1. As the prompt length increases, WER tends to decrease. When the length exceeds 16, the performance gain becomes less prominent. To strike a balance between performance and parameter efficiency, we opt to set the prompt length to 16 for all subsequent experiments. Table 1 also shows how the three reparameterization schemes influence the performance. A shared reparameterization MLP is even worse than the scheme without reparameterization. This could be attributed to that the shared MLP restricts the learning of inter-layer variations. This is proved by the fact that adopting separate MLPs for each layer leads to performance enhancements. As a result, we opt to employ separate MLPs for the reparameterization of soft prompts in our subsequent experiments.
3.3.2 SOTA Comparison
We compare our prompting scheme with other SOTA methods in two aspects: performance (evaluated in terms of WER) and parameter efficiency (evaluated in terms of task parameters). The results are shown in Table 2. While the prior works based on WavLM and TS-HuBERT have slightly better performance, the training is expensive and requires extensive structure modifications. Deep prompting, in contrast, is a much cheaper choice to extend Whisper for TS-ASR. We also report the results of LoRA, which is another PEFT method for adapting large models. It achieves the best performance, as shown in Table 2. However, despite performance, prompting methods have their unique advantages (e.g., supports multi-task parallel inference [14]). In deep prompting, employing reparameterization MLP gives an additional performance improvement on test-clean at the expense of an increase in the number of parameters to tune. However, once the training is completed, the MLP can be simply discarded. In this way, the number of task-specific parameters to be stored for inference remains limited.
@@ 3.3.2 SOTA Comparison
We then report the results of full fine-tuning in the last two rows of Table 2. Due to computing resource constraints, we are unable to provide the results for Whisper-large. A noticeable trend emerges when comparing full fine-tuning with the prompting approach: despite a substantial increase in the number of tunable parameters, full fine-tuning fails to exhibit satisfactory performance. The underlying cause of this performance gap is attributed to the training set train-100-clean is relatively small, causing severe over-fitting during full fine-tuning. This stark contrast underscores the effectiveness of prompt tuning in mitigating overfitting risks, especially when adaptation data for large-scale models are limited.
@@ 3.3.3 Ablation Study
We conduct ablation experiments to verify the impact of design components, including prompt tuning, deep prompting, and reparameterization MLP. As shown in Table 3, Whisper without any modification (i.e. zero-shot configuration) cannot perform well on target-speaker ASR. When we insert target speaker embedding together with soft prompts into both encoder and decoder input, a dramatic WER reduction, e.g. from 78.92% to 17.59% for Whisper-large on test-clean, is observed. This shows the efficacy of prompt tuning. By further adding the reparameterization MLP module, we observe performance improvement in most of the cases, showing that reparameterization of soft prompts can benefit the training. In addition, configurations with deep prompting show consistent performance improvement. This shows that more tunable parameters bring better performance and the soft prompts closer to the output layer have a more critical performance effect.
We further investigate how prompting location affects performance. Specifically, based on the best configuration (i.e., Whisper + PT + DP + MLP), we ablate the soft prompts inserted in the encoder/decoder to test the influence of varying prompting locations. Hereby, we set the prompt length to 32 to keep the number of parameters constant. The results are shown in the last two rows of Table 3, from which we see that the soft prompts inserted in the encoder play an important role in guiding the model while those inserted in the decoder play a relatively minor role but still lead to performance improvement. So, inserting soft prompts both in the encoder and decoder will lead to the best performance with a constant number of parameters.
@@ 3.3.4 Retaining Whisper’s Featured Abilities
We anticipate a practical TS-ASR system to not only accurately transcribe the desired speech of a target speaker but also to execute the entire speech processing pipeline. This includes various post-processing tasks, such as inverse text normalization and timestamp tagging, among others. The original Whisper has already demonstrated commendable performance in these aspects. However, it’s important to note that the prompted Whisper can no longer showcase its automatic punctuating and capitalizing abilities while performing TS-ASR. This limitation arises because the ground truth texts in LibriSpeech, which are manually labeled, lack case sensitivity and punctuation. Soft prompts have learned and perpetuated this bias. To retain the automatic inverse text normalization ability of Whisper, we propose to employ the auto-labeled texts generated by the original Whisper model on single-talker utterances as ground truth to supervise the training of soft prompts.
@@
The above table shows the results with and without inverse text normalization, where we highlight the automatically inserted punctuation and capitalized letters. We see that through training soft prompts supervised by the auto-labeled texts that are inverse-text-normalized, the prompted Whisper can now automatically perform inverse-text normalization along with target-speaker ASR with virtually no performance loss. By further removing the special token $\langle$|⟨ |no-timestamps|$\rangle$| ⟩ in multitask format (see Sec. 2.1), the prompted Whisper can produce reliable timestamps, as indicated by the highlights in the last row of the table, even though timestamps were not included in the adaptation data.
@@ 4 Conclusion
In this work, we have proposed a prompting scheme, that jointly leverages prompt tuning, and deep prompting together with the reparameterization technique, to extend the single-talker ASR model to the target-speaker ASR of multi-talker overlapped utterances. Our proposed method has shown comparable performance with SOTA and such performance was achieved with an attractively small amount of task parameters compared to existing methods. Notably, our method also retained the automatic inverse text normalization and timestamp prediction capabilities of the original Whisper model. This ensures that the generated transcriptions remain both natural and informative, further enhancing the versatility and utility of our method.